Pre-lab Assignment Experiment 21-A Simple Electrical Circuits

1. (2 pts) Capacitors need to be treated with a little bit of respect and used carefully. What are two things mentioned in the lab about which you should be careful when using capacitors?

2. Suppose a light bulb is rated for a voltage of 1.8 V. At this voltage it draws a current of 0.054 A.

a. (1 pt) Using Ohm’s Law, determine the resistance of the bulb. Show your work.

b. (2 pt) Draw a circuit that includes a battery and two of these bulbs in parallel.

c. (1 pt) What is the equivalent resistance of the two bulbs in parallel? Show your work.
Answer #1

1. (i) Do not charge by higher current or higher voltage than specified.

(ii) Do not reverse placement of (+) and (-).

2. Given V=1.8V and I = 0.054A

(a) From ohm's law

V=IR where R is resistance

therefore R=V/I = 1.8/0.054 =33.33 ohm

hope this will help you. Please give a thumbs up. thanks.